A high fever can interfere with normal neuronal activity. Since temperature is one of the terms in this equation (membrane potential = RT/ZF ln ([ion]in/[ion]out)) [R = 8.3145 J/K·mol; F = 96,485 J/V·mol], which defines membrane potential, a fever could possibly alter a neuron’s resting potential.
1. Calculate the effect of a change in temperature from 98.6°F to 105.8°F (37°C to 41°C) on a neuron’s membrane potential. Assume that the normal resting potential is -70mV and that the distribution of ions does not change. 2. How else might an elevated temperature affect neuronal activity?
Answer #1

If T = 98.6 °F = 37 °C = (37+273) K = 310 K

If T = 105.8 °F = 41 °C = (41+273) K = 314 K

Therefore, membrane potential at 98.6 °F = -0.07 V - 310R/ZF ln([ion]in/[ion]out)

And the membrane potential at 105.8 °F = -0.07 V - 314R/ZF ln([ion]in/[ion]out)

Hence, the change (increase) in temperature decreases the membrane potential.

2) And the elevated temperature decreases the neuronal activity.